# CoverUp: Coverage-Guided LLM-Based Test Generation Juan Altmayer Pizzorno

### v1

@@
Abstract
This paper presents CoverUp, a novel system that drives the
of coverage analysis and large-language models (LLMs). CoverUp
iteratively improves coverage, interleaving coverage analysis with
dialogs with the LLM to focus its attention on as yet uncovered
lines and branches. The resulting test suites significantly improve
coverage over the current state of the art:
compared to CodaMosa, a hybrid LLM / search-based software
testing system, CoverUp substantially improves coverage across the board.
On a per-module basis, CoverUp achieves median line coverage
of 81% (vs. 62%), branch coverage of 53% (vs. 35%) and
line+branch coverage of 78% (vs. 55%).
We show that CoverUp’s iterative, coverage-guided approach is crucial to its effectiveness, contributing to nearly half of its successes.
††footnotetext: Work done at the University of Massachusetts Amherst.
I Introduction
Test generation tools aim to increase the coverage of the program under test.
By generating tests that cover a wider range of scenarios and execution paths,
@@ I Introduction
Unfortunately, current approaches often fail to achieve substantial coverage.
Pynguin exemplifies the search-based software testing (SBST)
approach [1].
Starting from randomly created test cases, SBST employs genetic
algorithms to mutate the tests in an attempt to increase coverage.
Unfortunately, its search process can get
stuck. CodaMosa addresses this problem by tracking
Pynguin’s progress [2]; when it concludes that the search process has
stalled, it prompts a large language model (LLM) for a test. It then
uses that test to re-seed the SBST, allowing it to resume progress.
While CodaMosa greatly improves Pynguin’s performance and constitutes
the previous state of the art, the resulting code coverage can still remain
relatively low (e.g., 34% overall branch coverage;
see Section IV for details).
This paper proposes CoverUp, a novel approach that drives LLM-based test
generation with coverage analysis. CoverUp incorporates detailed
@@
To our knowledge, CoverUp is the first test generation system
to prompt an LLM for tests specifying the lines and branches lacking coverage
and that iteratively refines that prompt using coverage analysis.
We evaluate CoverUp using CodaMosa as a baseline, and on the
same benchmark applications originally used to evaluate it. Using the
state-of-the-art OpenAI’s GPT-4 Turbo LLM for both, we find that
CoverUp yields higher coverage in 61% of the benchmarks,
significantly increasing all coverage
metrics, whether computed across all code or aggregated over the
various benchmarks.
CoverUp increases median line coverage from 62% to 81%,
median branch coverage from 35% to 53% and
median line+branch coverage from 55% to 78%.
Across all code, CoverUp increases
line coverage from 54% to 61%, branch coverage from 34% to 43%,
and line+branch coverage from 49% to 57%. We show that CoverUp’s iterative, coverage-guided approach
is key to its effectiveness, accounting for roughly half of its successes.
@@
It presents CoverUp, a novel system that drives the
@@
It conducts an empirical analysis of CoverUp, showing that it
@@ II Related Work
specification-based [3],
random [4, 5],
feedback-directed random [6],
symbolic execution guided random (“concolic”) [7, 8, 9],
search-based software testing (SBST) [10, 11, 1, 12, 13]
and transformer based [14] approaches.
@@
Wang et al. survey 102 recent papers using LLMs for software testing [15].
@@
function under test’s implementation, its documentation and usage snippets [16].
refine the prompt in case the new tests do not improve coverage.
@@
Java test case generation, among other code generation tasks [17].
@@
As Section IV-D shows, this iterative dialogue is
@@
to generate property-based tests for Python [18, 19].
@@
a natural language description of the intended functionality of code [20].
@@
Fuzz4All [21] uses two separate LLMs to fuzz test programs in various programming
@@
CodaMosa [2], mentioned in the introduction, prompts
the Codex LLM for a test when its SBST algorithm gets stuck. Like
CoverUp, CodaMosa selects code to prompt for based on its (lack
of) coverage, and the prompt contains source code. However, it does
not use coverage information within the prompt itself, nor does it
systematically prompt for tests for all code lacking coverage; it only
prompts the LLM when necessary to help SBST make progress.
CodaMosa represents the previous state of the art in test case generation
for Python, and Section IV uses it as the baseline
for comparison to CoverUp.
Figure 1: Overview: The execution steps of CoverUp.
@@
III Technique
CoverUp first measures the code coverage of the existing test suite
and splits up the source code into segments that require more testing
(Section III-A).
It then prompts the LLM for tests for each of these segments, using a chat
interface
(Section III-B).
As the LLM generates tests, CoverUp executes them, again measuring coverage.
If the test does not compile, fails or lacks coverage, it continues the chat,
pointing out any problems and requesting improvements
(Section III-C).
Finally, having processed all code segments, CoverUp checks the extended
test suite, looking for any integration problems
(Section III-D).
Figure 1 provides a graphical overview of these steps.
III-ACode Segmentation
CoverUp’s first step is to measure the code coverage of any pre-existing test suite;
it uses SlipCover [22], a recently introduced coverage analyzer that has
near-zero overhead, for that task.
It then uses the AST of each source file missing coverage to identify code
segments that can be used for prompting.
The goal is to be able to provide to the LLM, with each prompt, a code excerpt
@@ III Technique
Keeping code segments short is important for several different reasons.
Older OpenAI LLMs such as GPT-3.5 and GPT-4, as well as on-premises LLMs like
Meta’s LLaMA have small input windows of up to a few
thousand tokens, limiting the maximum size of the code segment.
OpenAI’s most recent GPT-4 Turbo models allow for inputs of up to 128K tokens,
but this limit applies not only to the initial prompt, but to the total number of tokens
in the entire sequence of prompts and responses in the case of a continued chat.
Finally, OpenAI charges at a per-token rate in both prompts and responses,
so that it literally pays to be succinct.
Each code segment typically includes an entire function or method and all its contents.
If a segment contains a method, CoverUp also includes a few lines of the original
and definitions; see Figure 2 for an example.
In an effort to enhance code comprehension, CoverUp retains all comments in a
segment, although removing them could make the segment more concise.
@@
This limit is intended to guide CoverUp towards acceptable excerpt sizes,
facilitating its use with language models with more limited input window sizes.
@@
Figure 2: Example of a code segment with a class method:CoverUp created a code segment for a class method,
omitting over 200 lines of source code originally
@@
III-BInitial Prompting
CoverUp prompts the LLM for tests for every code segment it identifies;
Figure 3 shows an example of an initial prompt.
It has the following structure (circled numbers identify sections in the example):
a statement assigning the LLM the persona [23]
@@
a sentence pointing out the code excerpt (segment), identifying what file
and module it comes from, and stating what lines or branches do not execute
when tested.
@@
a request for a pytest test;
a series of other requests, such as “make sure that the new test is correct”
and “include assertions” to steer the result towards usable tests;
@@
the code segment with any relevant lines highlighted by prefixing
them with their line numbers.
Relevant lines are those that lack coverage or that are
part of a branch that lacks coverage.
23456Figure 3: Example of an initial prompt:
@@
and shows the code segment (6).
Any lines relevant to the request, i.e., used to identify
missing coverage, are highlighted by prefixing them with their line numbers
(as in “19:” in part (6)).
The code shown is from the flutils package.
Figure 4: Excerpt of test calling to pytest.main:
We found that certain tests generated by GPT-4
included top-level, unconditional calls into pytest.main.
Such calls, executed when pytest loads the test file,
caused it to restart the search for all tests in the suite,
slowing it down immensely.
Part (4) of the prompt now requests omitting such calls.
We found that tests generated by GPT-4 often included top-level code calling
@@
For that reason, part 4 of the prompt includes instructions that the test not
include such calls.
Prompt part 4 also instructs that the tests “clean up […], so as not
to affect other tests”.
It also suggests the use of pytest-mock, a library useful for isolating
the software under test from other parts of the code, facilitating test cleanup.
Despite these instructions, we still observe test generations that include side effects;
Section III-D describes how CoverUp detects and handles
such tests.
As Figure 3 shows, in part 6 of the prompt, CoverUp
prefixes certain lines with their numbers, such as in 20: before line 20.
It adds such prefixes to all lines lacking coverage as well as to all lines
@@
This improves the LLM’s understanding of the missing coverage:
we found that prompts that only indicated the line number at the beginning of the
excerpt led to more tests being generated that did not improve coverage.
If a code segment has no prior coverage, that is, every one of its lines lacks coverage,
then CoverUp uses a slightly modified initial prompt where part 2 simply
states “the code does not execute”.
Typically, the LLM will not generate enough tests to immediately yield high coverage
in response to this simpler prompt, but this can be improved upon with a subsequent
execution of CoverUp, where it refocuses the LLM on those lines missed in the first
iteration.
III-CVerification and Continued Chat
@@
again measuring coverage using SlipCover.
This step is made practical by SlipCover’s near-zero overhead:
@@
introduces a median of 180% overhead [22].
@@
But, if they do not increase coverage, or if they result in failures or errors,
@@
Figures 5 and 6 show examples
of those prompts.
12Figure 5: Example of a coverage follow-up prompt:CoverUp indicates to the LLM that a line and
@@
123Figure 6: Example of an error follow-up prompt:CoverUp indicates to the LLM that an error occurred (1),
@@
These are typically test helper modules, such as the pytest-ansible plugin
used to help test the ansible package, but could also be indicative of
a module required by the program under test that has not yet been installed.
CoverUp offers options to record the missing modules in a Python requirements.txt
format to facilitate making them available for a subsequent run.
III-DIntegration Check
@@
side effects (“state pollution”) that cause other tests to fail [24].
This issue is not unique to LLM-based test generation: Fraser et al. [25]
report side effects on 50% of test classes generated by EvoSuite [11],
@@
To attempt to detect and handle such tests, once CoverUp has processed all
outstanding code segments, it executes the entire suite in a single session.
If tests then fail, CoverUp attempts to identify the cause, disabling tests as
necessary by renaming their files so that pytest ignores them.
Such tests remain available for review and possible reactivation by the
user, as they are known to contribute to coverage.
CoverUp supports two different methods for handling failing tests.
It can simply disable all failing tests without attempting to ascertain
the cause of the errors.
This method is fast and allows the user to move on quickly to prompting
for more tests, but may lead to a significant loss in coverage as various
tests are disabled unnecessarily.
Alternatively, CoverUp can look for the test or tests actually causing the
failures by successively running subsets of the entire suite, an approach
adapted from Delta Debugging [26].
This method requires additional test runs, and thus more time, but is likely
to reduce the loss in coverage resulting from disabled tests.
Figure 7: Excerpt of a test with side effects:
@@
CoverUp detects these failures during its integration check and
optionally either disables all failing tests, or searches for the
test causing the failures.
III-EHandling Flaky Tests

### v2

@@
Abstract.
Testing is an essential part of software development.
Test generation tools attempt to automate the otherwise labor-intensive task of test creation,
but generating high-coverage
tests remains a challenge.
This paper proposes CoverUp, a novel approach to driving the
generation of high-coverage Python regression tests. CoverUp
iteratively improves test coverage, interleaving coverage analysis with
dialogs with the LLM that steer it to refine tests so that they increase coverage of
lines and branches.
We evaluate our prototype CoverUp implementation
across a benchmark of challenging
code derived from open-source Python projects, and show that
CoverUp substantially improves on the state of the art.
Compared to CodaMosa, a hybrid search/LLM-based test generator,
CoverUp achieves a per-module median line+branch
coverage of 80% (vs. 47%).
Compared to MuTAP, a mutation/LLM-based test generator,
CoverUp achieves an overall line+branch coverage of 90% (vs. 77%).
We show that CoverUp’s iterative, coverage-guided approach is
crucial to its effectiveness, contributing to nearly 40% of its successes.
Test Generation, Regression Testing, Large Language Models, Code Coverage
††ccs: Software and its engineering Software testing and debugging††footnotetext: Work done at the University of Massachusetts Amherst.
1. Introduction
Testing is essential to ensuring the quality of software, but
manually crafting tests can be so labor intensive that developers
choose not write them (Daka and Fraser, 2014).
Test generation tools attempt to automate this task, generally working
under the assumption that the source code is correct.
By generating tests based on it, they enable regression testing,
with the goal to prevent future bugs as the software is modified.
As these tools add tests that cover wider range of scenarios and execution paths,
@@ 1. Introduction
This paper proposes CoverUp, a novel approach to test generation
aimed aimed at achieving high coverage.
Our key insight is that large language models (LLMs) are capable of
simultaneously reasoning about code and coverage information.
We leverage this insight in the design of CoverUp.
CoverUp incorporates detailed
@@
Our empirical evaluation compares CoverUp to two state of the art
test generation systems: CodaMosa (Lemieux et al., 2023), a hybrid
search/LLM-based test generator, and MuTAP (Dakhel et al., 2024), a
a mutation/LLM-based approach.
We show that CoverUp substantially improves upon the state of the art.
Across a benchmark of challenging code derived from
open-source Python projects and using OpenAI’s GPT-4o LLM,
CoverUp increases coverage on every metric vs. CodaMosa, achieving
per-module median line+branch coverage of 80% (vs. 47%) and
overall line+branch coverage of 60% (vs. 45%).
Compared to MuTAP on one of the benchmark suites it supports,
CoverUp again achieves greater or equal coverage on every metric,
with an overall line+branch coverage of 90% (vs. 77%).
@@
It presents CoverUp, a novel approach that drives the
@@
It conducts a empirical analysis of our CoverUp prototype, showing that it
@@ 2. Related Work
specification-based (Boyapati et al., 2002),
random (Miller et al., 1990; Fioraldi et al., 2023),
feedback-directed random (Pacheco et al., 2007),
symbolic execution guided random (“concolic”) (Godefroid et al., 2005; Sen et al., 2005; Tillmann and de Halleux, 2008),
search-based software testing (SBST) (Fraser and Zeller, 2012; Fraser and Arcuri, 2011; Lukasczyk and Fraser, 2022; Panichella et al., 2015, 2018)
and transformer-based approaches (Tufano et al., 2020).
@@
Wang et al. survey 102 recent papers using LLMs for software testing (Wang et al., 2024).
@@
SBST approaches
Pynguin (Lukasczyk and Fraser, 2022) employs a search-based software testing (SBST)
approach. Starting from randomly created test cases, SBST employs genetic
algorithms to mutate the tests in an attempt to increase coverage.
Unfortunately, its search process can get stuck as test mutations repeatedly
lead to the same execution paths.
CodaMosa (Lemieux et al., 2023) addresses this problem by tracking
Pynguin’s progress; when it concludes that the search process has
stalled, it prompts an LLM for a test. It then uses that test to
re-seed the SBST, allowing it to resume progress.
We empirically compare CoverUp to CodaMosa in Section 4.2.
LLM-based Approaches
@@
Java test case generation, among other code generation tasks (Bareiß et al., 2022).
@@
As Section 4.4 shows, this iterative dialogue is
@@
to generate property-based tests for Python (Claessen and Hughes, 2000; Vikram et al., 2023).
@@
a natural language description of the intended functionality of code (Lahiri et al., 2022).
@@
function under test’s implementation, its documentation and usage snippets (Schäfer et al., 2024).
continue the chat if the new tests do not improve coverage, nor does it use tool functions
to provide the LLM with additional context.
ChatUniTest (Chen et al., 2024) prompts LLMs to generate Java unit tests.
Its use of LLMs is limited to prompting with code and for repairs when
a generated test fails compilation.
Unlike CoverUp, ChatUniTest does not employ coverage measurements
to indicate what lines or branches lack coverage, nor does it request
improvements if the generated tests do not improve coverage.
Concurrent Work Using LLM-based Approaches
The approaches below were developed concurrently with CoverUp, which was initially
posted on GitHub on August 7, 2023:
Fuzz4All (Xia et al., 2024) uses two separate LLMs to fuzz test programs in various programming
@@
SymPrompt (Ryan et al., 2024) extracts approximate path constraints from Python functions
through static analysis and prompts an LLM for tests based on those constraints.
In contrast, CoverUp prompts an LLM directly indicating the lines and branches that
lack coverage.
Whereas SymPrompt extracts related type declarations and other code and include
these statically in its prompts, CoverUp provides a tool function (OpenAI, 2023)
that looks up arbitrary additional information about types, functions and other symbols,
thus allowing the LLM to drive this process.
SymPrompt’s error handling is limited to deleting lines in the response in an attempt
to correct syntax errors; CoverUp continues the chat in case of errors, asking the
LLM for a correction.
CoverUp also continues the chat if the test does not improve coverage.
SymPrompt is evaluated on an undisclosed subset of 897 functions (from at least 4,546)
previously used by CodaMosa and BugsInPy (Widyasari et al., 2020).
It achieves 74% line coverage using OpenAI’s GPT-4 LLM, approximately
a 2x improvement over its baseline prompt (Ryan et al., 2024).
Without knowing the exact functions employed despite multiple requests for
code and/or a replication package111Personal communication with SymPrompt authors.,
we can only provide a coarse comparison:
as Section 4.2 shows, using GPT-4o CoverUp
obtains 82% median per-module line coverage and 64% overall line coverage
on suite CM.
The code in that suite was also derived from CodaMosa’s evaluation, but contains
4,116 functions.
Even though CoverUp is evaluated on an 5x larger suite, the median per-module
coverage it achieves is also approximately a 2x improvement over an ablated version
of itself (Section 4.2).
TestGen-LLM (Alshahwan et al., 2024) is a test generation tool developed and deployed at
Meta to improve Kotlin unit tests.
TestGen-LLM prompts LLMs for tests that improve coverage, including in the prompt the
class under test and in some cases the existing test class.
It discards any tests that do not build, do not pass, or do not improve coverage, and
presents any remaining tests to a developer for approval during code review.
Like TestGen-LLM, CoverUp rejects tests that fail or lack coverage, but instead of
simply discarding them, it continues the chat with the LLM, asking for improvements.
CoverUp also indicates in its prompts what portions of the code lack coverage, and
allow the LLM to discover additional context through a tool function.
MuTAP (Dakhel et al., 2024) prompts an LLM
for Python unit tests based on mutation testing.
MuTAP first prompts for a test for a portion of code and, using it,
performs mutation testing.
It then prompts for new assertions for each surviving mutant, adding
these to the test.
Like CoverUp, MuTAP also prompts the LLM for a repair in case of syntax errors,
but, unlike CoverUp, not for other kinds of errors, lack of coverage,
nor does it include coverage information in its prompts.
Unfortunately, MuTAP currently lacks any ability to provide context for the code
under test and hardcodes the datasets used for evaluation in the code,
hampering efforts to evaluate it on other code.
In contrast, our implementation of CoverUp, like Pynguin and CodaMosa,
can be applied to any Python package.
We empirically compare CoverUp to MuTAP in Section 4.2.
Figure 1. Graphical Overview of the CoverUp Algorithm.
@@
CoverUp provides a tool function with the prompts with which
the LLM can request additional context.
@@
3. Approach
This section presents a detailed description of CoverUp.
Figure 1 provides a graphical overview of its algorithm.
CoverUp first measures the code coverage obtained by the existing test suite and
uses it to identify code segments (functions and/or classes) that require additional testing
(Section 3.1).
It then prompts an LLM for tests for each of these segments,
combining information from coverage analysis and static code analysis
(Section 3.2).
CoverUp execute the LLM generated tests, once again measuring coverage.
If the tests do not compile, fail to run or increase coverage, CoverUp continues the
chat with prompts that request improvements and include any error messages.
(Section 3.3).
In both initial and continued prompts, CoverUp offers the LLM a tool function with
which it can request additional contextual information about names in the source code,
such as function or type definitions
(Section 3.4).
Having handled all code segments, CoverUp checks the extended test suite,
looking for any integration problems
(Section 3.5).
The entire process can be repeated an arbitrary number of times to allow the LLM
to re-focus on segments not covered by previous iterations.
3.1. Code Segmentation
CoverUp’s first step is to measure the code coverage of any pre-existing test suite.
It then uses the AST of each source file missing coverage to identify code segments
that need additional coverage.
Code segments typically consist of a single function or method; subject to a size limit,
however, they may contain an entire class.
As Figure 2 shows, if a segment contains a method,
CoverUp also includes a few lines of the original
and definitions.
Since we find that retaining comments improves code comprehension, CoverUp
does not remove these, thus trading a loss in concision for an increase in generation quality.
The goal is to provide to the LLM, with each prompt, a code excerpt
@@ 3.1. Code Segmentation
Keeping code segments short is important primarily because of LLMs’ limits on
their context (input) window.
While various current LLMs such as OpenAI’s GPT-4 and GPT-4o, Meta’s LLaMA 3 and
Anthropic’s Claude 3 support large context windows, of over 100K tokens,
this limit applies not only to the initial prompt, but to the total number of tokens
in the entire sequence of prompts and responses in the case of tool function calls
or a continued chat.
Even when prompts fit in these long context windows, shorter prompts remain preferable:
Rosas et al. (Rosas et al., 2024) show that
as the size of prompts containing code increases, so does the likelihood of inaccuracies.
Finally, LLM API use is often charged at a per-token rate, so that it literally
pays to be succinct.
@@
CoverUp also generates import statements that provide context
for each code segment: we find that without these, LLMs often make incorrect
assumptions about symbols in the code, leading to errors.
Rather than extract these verbatim from source code, CoverUp generates imports
based on static analysis.
In particular, it converts any relative imports to absolute imports: using relative
imports in code excerpts leads to these being used in tests outside of package
scope, where they are not valid.
We implement this step using SlipCover (Altmayer Pizzorno and Berger, 2023), a recently introduced
coverage analyzer that has near-zero overhead, and Python’s ast module.
Functionidentify_segments(coverage, max_len)
foreachindo
lines_in(missing_branches(coverage, file))
foreachindo
empty list
while( is-a Class) andanddo
append to
return
Algorithm 1Algorithm for identifying the code segments lacking coverage, based on a coverage measurement and a tentative maximum segment length.
Figure 2. CoverUp summarizes method excerpts:CoverUp generates extremely compact code excerpts.
In this case, 200 lines of source code were originally
@@
3.2. Initial Prompting
CoverUp next prompts the LLM for tests for each code segment it identified in the previous step.
Figure 3 shows an example of an initial prompt, with circled numbers
identifying sections.
It has the following structure:
a statement assigning the LLM the persona (White et al., 2023)
@@ 3.2. Initial Prompting
a sentence pointing out the code excerpt (segment), identifying what
file it comes from, stating what lines or branches do not execute when tested.
@@
a request for a pytest test functions, and an
encouragement for the LLM to use the provided tool function;
a series of other requests, such as
“include assertions” and “avoid state pollution”, to steer the result towards usable tests;
@@
the code segment, prefixed by generated import statements
and tagging the lines lacking (line or branch) coverage with their numbers.
[…]
@@
and shows the code segment, prefixed by import statements and
with the line lacking coverage identified by its number.
The code shown is from the thonny package.
We find that tests generated by GPT-4 often include top-level code calling
@@
For that reason, part 4 of the prompt directs the LLM to
not include such calls.
Prompt part 4 also instructs the LLM to generate tests that
“clean up […], so as to avoid state pollution”.
It suggests the use of monkeypatch and pytest-mock, useful for
automatically cleaning up after tests and isolating the software under test from other parts of the code.
Including these instructions improves the generated tests; nonetheless
we still observe test generation that includes side effects.
Section 3.5 describes how CoverUp handles such tests.
Within the excerpt in part 6, CoverUp tags certain lines with
their numbers, as in in “19:” in Figure 3.
It tags all lines lacking coverage as well as all lines
@@
This tagging improves the LLM’s understanding of the missing coverage:
we find that prompts that only indicate the starting line number at the beginning of the
excerpt lead to more tests being generated that do not improve coverage.
3.3. Verification and Continued Chat
@@
again measuring coverage.
In our implementation, this step is made more efficient by SlipCover’s near-zero overhead:
@@
introduces up 260% overhead (Altmayer Pizzorno and Berger, 2023).
@@
If they do not increase coverage, or if they result in failures or errors,
@@
Figures 4 and 5 show examples
of these prompts.
[…]
@@
[…]
@@
These missing modules are typically test helper modules, such as the pytest-ansible plugin
used to help test the ansible package;
Our implementation offers options to automatically install missing modules and to
record these in Python’s requirements.txt format to facilitate making
their use in a subsequent run.
3.4. Tool Functions
Tool functions allow an LLM to interact with external tools (OpenAI, 2023).
CoverUp exposes a get_info tool function that the LLM
can call to request additional information about any names in the excerpt,
such as functions, classes, variables or modules.
It responds to such function calls by performing static analysis and slicing relevant
portions of the source code.
Figure 6. Example of a prompt:
In response to this initial prompt, shown abbreviated, the LLM might
respond with a get_info("code.A") function call, as requires
more information on how to instantiate A in tests.
Figure 7. Example function call response:
Asked about code.A, CoverUp would perform static analysis and,
discovering that A is a class, respond with a slice of its definition.
In this case, it includes its __init__ constructor, which indicates to
the LLM that A’s constructor takes the initial_x keyword
argument.
The slice does not include any of the contents of method m; the LLM could
subsequently ask for code.A.m to obtain more information on it.
For example, when given the prompt shown in Figure 6, the LLM
might ask for information on code.A.
The function response, shown in Figure 7, indicates that values
of x must be passed to the class constructor using a keyword argument.
This response lets the LLM immediately respond with correct tests; we find that if we do not
include the tool function in this situaiton, the LLM attempts to simply pass values of x to the
constructor, leading to errors.
CoverUp makes get_info available both with initial (Section 3.2),
and continuation prompts (Section 3.3), so that it can help generate
both better responses and better corrections.
3.5. Integration Check
@@
side effects (“state pollution”) that cause other tests to fail (Gyori et al., 2015).
This issue is not unique to LLM-based test generation: Fraser et al. (Fraser and Arcuri, 2014)
report side effects on 50% of test classes generated by EvoSuite (Fraser and Arcuri, 2011),
@@
For a test generation approach to be effective, it must thus include some way to
handle state pollution.
Our implementation offers three alternatives:
Execute each test in isolation:
This approach prevents memory-based state
pollution by one test from affecting other tests.
CoverUp implements efficient test isolation based on the Unix fork()
system call for the pytest framework.
This option is the default, which Section 4 uses
for the empirical evaluation of CoverUp.
Disable polluting tests:
This approach executes the entire suite and, upon a test failure,
searches for the polluting tests and disables them.
CoverUp implements the search by successively reducing and executing subsets
of the entire suite;
Disable failing tests:
This approach may lead to a significant loss in coverage as tests may
be disabled unnecessarily, and may still leave a polluting test enabled,
but allows the user to move on quickly to prompting for more tests.
Any disabled tests remain available for review and possible reactivation by the user.
Figure 8. Handling state pollution:
@@
Our implementation offers three alternatives to handle these
situation: it can execute tests in isolation (the default), so that
and disable the polluting tests, or disable the failing tests.
3.6. Handling Flaky Tests